### v4

# Self-supervised Learning of Hybrid Part-aware 3D Representations of 2D Gaussians and Superquadrics

Abstract
Low-level 3D representations, such as point clouds, meshes, NeRFs and 3D Gaussians, are commonly used for modeling 3D objects and scenes. However, cognitive studies indicate that human perception operates at higher levels and interprets 3D environments by decomposing them into meaningful structural parts, rather than low-level elements like points or voxels. Structured geometric decomposition enhances scene interpretability and facilitates downstream tasks requiring component-level manipulation. In this work, we introduce PartGS, a self-supervised part-aware reconstruction framework that integrates 2D Gaussians and superquadrics to parse objects and scenes into an interpretable decomposition, leveraging multi-view image inputs to uncover 3D structural information. Our method jointly optimizes superquadric meshes and Gaussians by coupling their parameters within a hybrid representation. On one hand, superquadrics enable the representation of a wide range of shape primitives, facilitating flexible and meaningful decompositions. On the other hand, 2D Gaussians capture detailed texture and geometric details, ensuring high-fidelity appearance and geometry reconstruction. Operating in a self-supervised manner, our approach demonstrates superior performance compared to state-of-the-art methods across extensive experiments on the DTU, ShapeNet, and real-world datasets.
footnotetext: Corresponding authors
1 Introduction
Figure 1: Prior works [32, 36] model the scene using primitives, while the proposed method can further model precise geometry details and textures. Here, EMS [32] takes point cloud inputs and reconstructs non-textured primitives.
3D reconstruction from multi-view images is a long-standing challenge in 3D vision and graphics [55, 17, 12]. Most reconstructed scenes are in low-level representations such as point clouds, voxels, or meshes. However, humans tend to understand 3D scenes as reasonable parts [35]. For instance, when observing a scene, we naturally construct high-level structural information, such as scene graphs, instead of focusing on low-level details like point clouds or voxels. Motivated by it, we propose a part-aware reconstruction framework that decomposes objects or scenes into meaningful shapes or parts, facilitating tasks such as physical simulation, editing, content generation and understanding [18, 28, 58, 53, 54].
Several prior works [72, 37, 29, 32, 33, 56] have explored part-aware reconstruction or 3D decomposition. However, most of them rely heavily on 3D supervision and often struggle to retain fine-grained geometric details, limiting their practicality in real-world scenarios. Recent advances [48, 68] have focused on extending the neural radiance field framework for part-aware reconstruction, building upon its remarkable success in 3D reconstruction from multi-view images. For example, PartNeRF [48] models objects using multiple neural radiance fields. Yet, the intricate composition of implicit fields complicates the learning process, leading to suboptimal rendering quality and inefficient decomposition. Recently, DBW [36] proposes a novel approach that decomposes scenes into block-based representations using superquadric primitives [2], optimizing both their parameters and UV texture maps through rendering loss minimization. While this method demonstrates effective scene decomposition into coherent components, it faces challenges in achieving accurate part-aware reconstruction of both geometry and appearance, as illustrated in Fig. 1.
To address these limitations, we propose PartGS, a self-supervised part-aware hybrid representation that integrates 2D Gaussians [21] and superquadrics [2], to achieve both high-quality texture reconstruction and geometrically accurate part decomposition. There are previous approaches [20, 14, 50] combining mesh reconstruction and Gaussian splatting, where they first reconstruct the mesh and then attach Gaussians to it. In contrast to their primary focus on acquiring the representations, the proposed approach is to achieve part-aware decomposition reconstruction. Our method establishes a coupled optimization framework that simultaneously learns superquadric meshes and Gaussians through parameter sharing, with each part constrained to a single superquadric. The differentiable rendering of Gaussians drives this hybrid representation, leveraging the inherent convexity of superquadrics for qualitative 3D decomposition. The self-supervised training of part-aware reconstruction is built upon the assumption that each part of most objects should be a basic shape and can be represented by a superquadric. Meanwhile, Gaussian splatting, renowned for its superior rendering quality and efficient training, is incorporated to capture intricate texture details, speeding up reconstruction and improving rendering quality.
In the hybrid representation, 2D Gaussians are distributed on superquadric surfaces to form structured blocks. The pose and shape of Gaussians within each block are determined by their corresponding superquadrics rather than being optimized independently. The parameter space encompasses global controls for superquadric properties (shape, pose, and opacity) and local spherical harmonic coefficients for individual Gaussians. Compared to standard Gaussian splatting [25, 21], which populates the occupancy volume with independently parameterized Gaussians, the coupled representation is more compact and efficient.
During training, the parameters are optimized through rendering loss without additional supervision. However, we notice that image rendering loss often leads to local minima in superquadric shape optimization. To tackle this, we introduce several regularizers to maintain global consistency between the 3D representation and the input 2D information. This strategic implementation allows us to segment parts fully self-supervised, achieving block-level reconstruction. Finally, to better capture irregular shapes, we implement a point-level refinement step that frees 2D Gaussians to deviate from the surface, thereby enhancing geometric fidelity. Extensive experiments show that PartGS, at block-level and point-level, achieves 33.3% and 75.9% improvements in reconstruction accuracy, 3.18 and 16.13 increases in PSNR, and 4X and 3X speedups compared to the state-of-the-art baseline. Our contributions are summarized as follows:
  • We introduce a novel hybrid representation for part-aware 3D reconstruction, combining the strengths of superquadrics and Gaussian splatting to achieve reasonable decomposition and high-quality rendering.
  • We propose several novel regularizers to enforce consistency between 3D decomposition and 2D observations, enabling self-supervised part decomposition.
  • Compared to prior works, the method takes one step forward, by reconstructing both the block-level and point-level part-aware reconstructions, preserving both part segmentation and reconstruction precision.
Figure 2: Overview of our pipeline. PartGS takes multi-view images to learn a parametric hybrid representation of superquadrics and 2D Gaussians. It initializes from random superquadrics and is gradually optimized during training to obtain a block-level reconstruction. Then, we free the constraints of Gaussians to model detailed geometry, achieving point-level reconstruction.
2 Related Work
2.1 Shape Decomposition and Abstraction
Structured shape representation learning decomposes objects or scenes into coherent geometric primitives, facilitating shape understanding and generation [59, 19, 15, 9, 47]. Early approaches like Blocks World [43] and Generalized Cylinders [3] emphasized compact representations. Modern methods typically process 3D inputs (point clouds, meshes, voxels) by decomposing them into primitive ensembles, including cuboids [49, 42], superquadrics [38, 39, 60, 32], and convex shapes [8, 10]. For instance, MonteBoxFinder [42] integrates clustering [11, 51, 62], cuboid representations, and Monte Carlo Tree Search, while EMS [32] adopts a probabilistic approach for superquadric recovery. However, these methods are limited to coarse shape representations. Recent advances in shape abstraction [13, 40, 22, 46, 31] enable more detailed shape representations through flexible primitive deformation.
Some studies [48, 36, 68, 1, 16, 70] attempt to create structure-aware 3D representations directly from images. PartNeRF [48] introduces ellipsoid representations within NeRFs, its reliance on multiple implicit fields results in inefficient 3D decomposition. ISCO [1] and DBW [36] use 3D superquadrics for shape decomposition, which enables more meaningful structural separation. However, their simple shape parameters struggle to capture complex geometries, leading to poor geometry and appearance reconstruction. DPA-Net [68] has advanced 3D shape abstraction from sparse views but generates redundant parts and struggles with realistic texture rendering. A concurrent work, GaussianBlock [24], employs SAM [27] to guide superquadric splitting and fusion for 3D decomposition, yet its computational efficiency remains limited, typically requiring several hours per scene. In contrast, our approach accomplishes self-supervised [61, 52] part-aware scene reconstruction through an efficient hybrid representation that simultaneously maintains geometric fidelity and photorealistic rendering quality.
2.2 Mesh-based Gaussian Splatting
Gaussian splitting (GS) [25] has been rapidly adopted in multiple fields due to its remarkable rendering capability. Several studies [20, 7, 14, 50] attempt to align Gaussians with mesh surfaces for easier editing and animation. Among them, SuGaR [20] uses flat 3D Gaussians to enforce the alignment with the scene surface during optimization, minimizing the difference between the signed distance functions (SDF) of the desired Gaussians and actual Gaussians. GaMeS [50] introduces a hybrid representation of Gaussians and mesh, where Gaussians are attached to triangular facets of the mesh. Similarly, Gao et al. [14] proposes a mesh-based GS to achieve large-scale deformation effects on objects. Recently, 2DGS [21] proposes 2D Gaussians for surface modeling and significantly enhances the geometric quality. Inspired by them, we propose a representation that combines 2D Gaussians with superquadrics. A key distinction between our method and previous mesh-based GS approaches is that these methods require first reconstructing the scene’s mesh and then binding Gaussians to the mesh surface, which results in a non-continuous mesh. In contrast, our method directly optimizes the mesh through a rendering loss, enabling part-aware mesh reconstruction.
3 Method
Given a set of calibrated multi-view images and foreground masks, we aim to learn part-aware 3D representations, a meaningful decomposition of both geometry and appearance. Our approach adopts a two-stage optimization strategy: first, decomposing the object into basic shapes using a mixture of Gaussians and superquadrics at the block-level, followed by refining the decomposition at the point-level to achieve precise geometry. In Sec 3.1, we parameterize the hybrid representation. Sec. 3.2 then elaborates on leveraging this representation, enhanced by novel regularizers and an adaptive control strategy, to achieve self-supervised block-level decomposition. Finally, Sec. 3.3 presents the process for obtaining detailed part-aware results.
3.1 Parametrizing the Hybrid Representation
As shown in Fig. 2, to leverage the strengths of both, we attach Gaussians to the surface of superquadric meshes. This representation retains the superquadric’s ability to parse a 3D scene into distinct parts. Meanwhile, spherical harmonics of Gaussians enable complex texture rendering via Gaussian splatting, addressing the texture learning limitations of prior work [32, 42, 36, 67]. Sharing pose parameters between superquadrics and 2D Gaussians further improves the representation’s efficiency.
The parametric representation is controlled by both primitive and Gaussian parameters, which are optimized simultaneously. Given a 3D scene $\mathcal{S}$, the proposed method decomposes it into multiple hybrid blocks, each consisting of a superquadric with associated Gaussians. Each scene is denoted by the hybrid representation: $\mathcal{S}=\mathcal{B}_{1}\cup...\mathcal{B}_{i}...\cup\mathcal{B}_{M}$, where $\mathcal{B}_{i}$ is the $i$-th hybrid block, and $M$ is the total number of blocks. Blocks are defined using manually designed parameters that control pose, opacity, scale, shape, and texture. These parameters are optimized via differentiable rendering to parse the target scene.
Shape and Scale Parameters. For each hybrid block $\mathcal{B}_{i}$, its geometry is controlled by superquadric parameters [2]. Specifically, there are two shape parameters $\epsilon_{1},\epsilon_{2}$ that define its shape, along with three scale parameters $s_{1},s_{2},s_{3}$, which scale the 3D axes. Analogous to icosphere vertex positioning, superquadric vertex coordinates are computed by:
$\footnotesize\mathbf{v}=[s_{1}\cos^{\epsilon_{1}}(\theta)\cos^{\epsilon_{2}}(\varphi);s_{2}\sin^{\epsilon_{1}}(\theta);s_{3}\cos^{\epsilon_{1}}(\theta)\sin^{\epsilon_{2}}(\varphi)],$ (1)
where $\theta$ and $\varphi$ represent the azimuth and elevation defined in the spherical coordinate. The shape and scale parameters govern block deformation to learn part-aware geometry.
Pose Parameters. The pose of the $i$-th hybrid block is defined by its rotation $\mathbf{R}_{i}$ and translation $\mathbf{t}_{i}$. The vertex position from Eq. 1 is transformed from the local coordinate to world space as: $\hat{\mathbf{v}}^{j}_{i}=\mathbf{R}_{i}\mathbf{v}^{j}_{i}+\mathbf{t}_{i}$, where $j$ indexes the vertices. Previous approaches, including SuGaR [20] and GaMeS [50], position Gaussians on reconstructed meshes with independent pose and shape parameters to enhance appearance modeling. In contrast, our method employs Gaussians to construct a differentiable rendering that bridges images and superquadrics. To achieve this, we couple the parameters of Gaussians with superquadrics.
For a posed superquadric, Gaussian centers are uniformly sampled on triangular faces, with their poses determined by face vertices. Following GaMeS [50], each Gaussian’s rotation matrix $\mathrm{R}_{v}=[r_{1},r_{2},r_{3}]$ and scaling $\mathrm{S}_{v}$ are computed from vertex positions. Given a triangular face $V$ with vertices $v_{1},v_{2},v_{3}\in\mathbb{R}^{3}$, orthonormal vectors are constructed such that $r_{1}$ aligns with the face normal, $r_{2}$ points from the centroid $m=\text{mean}(v_{1},v_{2},v_{3})$ to $v_{1}$. $r_{3}$ is obtained by orthonormalizing the vector from the center to the second vertex concerning the existing $r_{1}$ and $r_{2}$ :
$r_{3}=\frac{{ort}(v_{2}-m;r_{1},r_{2})}{||{ort}(v_{2}-m;r_{1},r_{2}||},$ (2)
where ${ort}$ represents the orthonormalization in Gram-Schmidt [4]. For the scaling of 2D Gaussians, we use:
$\mathrm{S}_{v}=\text{diag}(s_{v}^{2},s_{v}^{3}),$ (3)
where $s_{v}^{2}=c||m-v_{1}||$, $s_{v}^{3}=c<v_{2},r_{3}>$, and $c$ is a size control hyperparameter. We place a fixed number of Gaussians on each triangular face. These designs eliminate the need to learn the geometry of Gaussians, thereby enhancing the efficiency of the representation.
Opacity Parameters. We define the total number of hybrid blocks as $M$. However, a typical scene does not contain exactly $M$ blocks. Therefore, only the meaningful blocks are retained. To achieve this, we introduce a learnable parameter $\tau_{i}$ to represent each block’s opacity. During the optimization, only blocks with $\tau_{i}$ greater than a certain threshold are retained. Note that Gaussians within the same block share the same $\tau$ for opacity in rasterization.
Texture Parameters. The texture is modeled using 2D Gaussians positioned on the surface of the superquadrics. Spherical harmonics of each Gaussian control the texture and are optimized for rendering view-dependent images [25].
3.2 Block-level Decomposition
This section describes the optimization of the hybrid representation. We observed that minimizing the rendering loss across multi-view images alone led to instability in positioning hybrid blocks. Therefore, several regularization terms are introduced to optimize the composition for maximal image formation consistency.
3.2.1 Optimization
In addition to the image rendering loss, we employ regularizers consisting of coverage, overlap, parsimony, and opacity entropy. The coverage loss encourages hybrid blocks to cover only meaningful regions; the overlap loss prevents blocks from overlapping; the parsimony loss regularizes the number of existing blocks; and the opacity entropy encourages block opacities close to binary.
Rendering Loss. The rendering loss is based on 3DGS [25], where it combines an $L_{1}$ term with a D-SSIM term:
${\mathcal{L}}_{\text{ren}}=(1-\lambda)\ L_{1}+\lambda L_{\text{D-SSIM}}.$ (4)
Coverage Loss. The coverage loss ensures that the block set $\{\mathcal{B}_{1}\dots\mathcal{B}_{M}\}$ covers the object while preventing it from extending beyond its boundaries. To determine the 3D occupancy field based on the blocks, we first define the approximate signed distance of a 3D point $x$ to the $i$-th block, $D_{i}(x)=\Psi_{i}(x)-1$. Here, $\Psi_{i}$ denotes the superquadric inside-outside function [2] associated with the block $i$. Consequently, $D_{i}(x)\leq 0$ if the point $x$ lies inside the $i$-th block, and $D_{i}(x)>0$ if $x$ is outside the block. Inspired by NeRF [34], we sample a ray set $\mathcal{R}$ using ray-casting based on camera poses. Given the object mask, we associate each ray $r\in\mathcal{R}$ with a binary label: $l_{r}=1$ if the ray $r$ is inside mask, otherwise $l_{r}=0$. The coverage loss is defined as:
${\mathcal{L}}_{\text{cov}}({\mathcal{R}})=\sum_{r\in{\mathcal{R}}}l_{r}L_{\text{cross}}(r)+(1-l_{r})L_{\text{non-cross}}(r).$ (5)
Here, $L_{\text{cross}}$ encourages rays to intersect with blocks, while $L_{\text{non-cross}}$ discourages rays from intersecting with blocks:
$L_{\text{cross}}(r)=\text{ReLU}\Big{(}\min_{i\in\mathcal{M}}\min_{x_{j}^{r}\in{\mathcal{X}}_{r}}D_{i}(x_{j}^{r})\Big{)},$ (6)
$L_{\text{non-cross}}(r)=\text{ReLU}\Big{(}\max_{i\in\mathcal{M}}\max_{x_{j}^{r}\in{\mathcal{X}}_{r}}-D_{i}(x_{j}^{r})\Big{)},$ (7)
where $x_{j}^{r}$ means the $j$-th sampled point along the ray $r$ and =$\mathcal{M}=\{1,\dots M\}$. Intuitively, this implies that at least one sampled points ${\mathcal{X}}_{r}$ along the ray $r$ inside the mask that lies within a certain block, while all points along the ray outside the mask should not belong to any block.
Overlap Loss. We introduce a regularization term to prevent overlap between individual blocks. Given the difficulty of directly calculating block overlap, we adopt a Monte Carlo method similar to  [36]. Specifically, we sample multiple 3D points in space and penalize those that lie inside more than $k$ blocks. Based on the superquadric inside-outside function, a soft occupancy function is defined as:
$\mathcal{O}_{i}^{x}=\tau_{i}\text{ sigmoid}(\frac{-D_{i}(x)}{\gamma}),$ (8)
where $\gamma$ is a temperature parameter. Thus, for a set of $N$ sampled 3D points $\Omega$, the overlap loss is expressed as:
${\mathcal{L}}_{\text{over}}=\frac{1}{N}\sum_{x\in\Omega}\text{ReLU}\left(\sum_{i\in\mathcal{M}}\mathcal{O}_{i}^{x}-k\right).$ (9)
Parsimony Loss. To promote the use of the minimal number of blocks and achieve parsimony in decomposition, we introduce a regularization term that penalizes block opacity ($\tau$). This loss is defined as:
${\mathcal{L}}_{\text{par}}=\frac{1}{M}\sum_{i\in\mathcal{M}}\sqrt{\tau_{i}}.$ (10)
Opacity Entropy Loss. During optimization, the opacity of the block inside the object region tends to approach 1, while the opacity of the block outside the object region tends to approach 0. To facilitate this, we associate the opacities of blocks with the labeled rays and apply a cross-entropy loss $L_{ce}$ between the block opacity and the mask labels, defined as follows:
${\mathcal{L}}_{\text{opa}}({\mathcal{R}})=\frac{1}{|{\mathcal{R}}|}\sum_{r\in{\mathcal{R}}}L_{ce}\left(\max_{i\in\mathcal{M}}\tau_{i}(x^{r}),l_{r}\right)~{}.$ (11)
Here, only points $x^{r}$ inside the blocks are sampled, e.g., $\left\{x^{r}\in{\mathcal{X}}_{r}\mid D_{i}(x^{r})\leq 0\right\}$.
The total loss is the weighted sum of the loss terms described above:
${\mathcal{L}}={\mathcal{L}}_{\text{ren}}+\lambda_{\text{cov}}{\mathcal{L}}_{\text{cov}}+\lambda_{\text{over}}{\mathcal{L}}_{\text{over}}+\lambda_{\text{par}}{\mathcal{L}}_{\text{par}}+\lambda_{\text{opa}}{\mathcal{L}}_{\text{opa}}.$ (12)
3.2.2 Adaptive Number of Blocks.
Given that the number of parts may vary across different scenes, we allow the number of blocks to adjust during optimization adaptively. Specifically, when the opacity of an existing block falls below a threshold $t$, this block is removed immediately. Furthermore, we explore a block-adding mechanism to dynamically incorporate new components, ensuring comprehensive coverage of target objects. Specifically, DBSCAN [45] is employed to cluster initial point clouds that are not covered by any blocks, where the point clouds are either randomly initialized or derived as by-products from COLMAP [44]. For each point cloud cluster, a new block is introduced at its center, with the remaining parameters of the block initialized randomly.
The final block-level reconstruction generates a textured 3D scene represented by multiple superquadrics, with surface details rendered through Gaussians, as illustrated in Fig. 2.
3.3 Point-level Decomposition
The primitive-based hybrid representation effectively decomposes the shape into parts but exhibits suboptimally for complex objects. To address this limitation, we perform a refinement stage that enhances geometric fitting. In this stage, the constraint between Gaussians and superquadrics is decoupled, enabling independent Gaussian optimization. Furthermore, to prevent Gaussian components from one block passing through adjacent blocks and disrupting the continuity and plausibility of the part decomposition, we impose a new constraint to minimize the negative signed distance of each Gaussian entering other blocks:
${\mathcal{L}}_{\text{enter}}=\frac{1}{N}\sum_{x\in\Omega}\sum_{m\in\mathcal{M}\setminus\{\delta\}}\text{ReLU}(-D_{m}(x)),$ (13)
where $\Omega$ is the sampled Gaussian set, and $\delta$ represents the identifier of the block to which the Gaussian $x$ belongs. As seen in Fig. 2, the reconstruction becomes more accurate and aligns better with the target shape.
Figure 3: Qualitative comparison on DTU [23] and ShapeNet [23] . The first two rows are DTU examples, and the last two are ShapeNet examples, respectively. Our method is the only one that provides reasonable 3D part decomposition while capturing detailed geometry.
Figure 4: Qualitative results on BlendedMVS [63] and self-captured data. We present RGB renderings and decomposed parts from novel views. The top examples are from the BlendedMVS dataset, and the last example is from our captured scenes.
Figure 5: Ablation studies on key strategies. The block-level visual comparisons illustrate the impact of adopting our proposed strategy. The first row shows results without the strategy, and the second with the strategy implemented.
Chamfer distance per scene Mean Mean
Method Input Renderable S24 S37 S40 S55 S63 S65 S69 S83 S97 S105 S106 S110 S114 S118 S122 CD #P
EMS [32] 3D GT 6.32 3.54 2.99 4.30 4.16 4.01 3.75 3.24 4.97 4.34 4.16 7.62 7.58 4.46 4.03 4.65 7.7
MBF [42] 3D GT 3.12 2.66 3.84 2.54 1.59 2.11 2.19 2.01 2.32 2.45 2.17 2.12 3.83 2.02 2.55 2.50 34.1
EMS [32] + Neus [55] Image 5.99 5.56 4.43 4.32 5.42 6.14 3.75 3.96 4.63 4.34 5.88 5.11 4.29 4.83 3.53 4.97 8.87
MBF [42] + Neus [55] Image 2.69 3.37 3.22 2.69 3.63 2.60 2.59 3.13 2.85 2.51 2.45 3.72 2.24 2.49 2.52 2.85 46.7
PartNeRF [48] Image 9.38 10.46 9.08 8.63 6.04 7.25 7.22 9.15 8.72 10.01 6.72 9.85 7.85 8.68 9.21 8.54 8.0
DBW [36] Image 5.41 8.35 1.57 3.08 3.40 4.15 7.46 3.94 6.63 4.85 4.38 4.65 6.29 4.34 3.04 4.76 4.8
Ours (Block-level) Image 5.68 4.91 1.85 2.61 3.75 4.66 3.75 7.57 4.27 4.38 3.49 4.48 3.61 4.21 3.70 4.19 5.9
Ours (Point-level) Image 0.70 1.17 0.55 0.65 1.06 1.23 1.10 1.36 1.37 0.78 0.92 1.41 0.69 1.05 0.71 0.98 5.9
Table 1: Quantitative comparison on DTU [23]. The Chamfer distance between the 3D reconstruction and the ground-truth is reported in 15 scenes. The best results are bolded, and the average numbers of primitives found (#P) that are greater than 10 are underlined.
Method Part-aware CD $\downarrow$ PSNR $\uparrow$ SSIM  $\uparrow$ LIPPS $\downarrow$ Time $\downarrow$
[…]
2DGS [21] 0.81 34.07 0.99 0.019 $\sim$ 10 m
PartNeRF [48] 9.59 17.97 0.77 0.246 $\sim$ 8 h
DBW [36] 4.73 16.44 0.75 0.201 $\sim$ 2 h
Ours (Block-level) 4.19 19.84 0.82 0.189 $\sim$ 30 m
Ours (Point-level) 0.98 35.04 0.99 0.015 $\sim$ 40 m
Table 2: Quantitative results on DTU [23]. Our method outperforms all part-aware approaches in image synthesis quality, reconstruction accuracy, and efficiency. Neuralangelo’s results are from the original paper, with all times measured on an RTX 3090 GPU.
Chamfer Distance  $\downarrow$ Primitives (#P ) Mean Mean
Method Input Airplane Table Chair Gun Airplane Table Chair Gun CD #P
EMS [32] 3D GT 3.40 6.92 19.0 2.02 9.4 7.88 10.3 8.4 7.84 9.0
MBF [42] 3D GT 2.83 2.18 1.59 2.32 10.85 13.9 13.4 14.3 2.21 13.1
PartNeRF [48] Image 2.29 2.77 2.30 2.46 8.0 8.0 8.0 8.0 2.46 8.0
DBW [36] Image 3.61 7.33 6.19 2.09 2.7 5.2 3.6 3.3 4.81 3.7
Ours (Block-level) Image 2.47 2.15 2.32 1.78 3.9 6.6 7.6 5.0 2.18 5.8
Ours (Point-level) Image 1.29 1.72 0.94 1.07 3.9 6.6 7.6 5.0 1.25 5.8
Table 3: Quantitative comparison on ShapeNet [6]. We report Chamfer distance and the number of parts. The best results are bolded, and the second-best results are underlined.
Chamfer Distance  $\downarrow$ PSNR  $\uparrow$ SSIM  $\uparrow$ LIPPS  $\downarrow$
Method Part-aware Airplane Table Chair Gun Airplane Table Chair Gun Airplane Table Chair Gun Airplane Table Chair Gun
2DGS [21] 1.47 2.37 0.50 1.03 40.89 39.80 39.05 41.74 0.994 0.990 0.990 0.994 0.009 0.027 0.017 0.009
PartNeRF [48] 2.29 2.77 2.30 2.46 19.63 20.66 19.08 21.97 0.898 0.855 0.875 0.916 0.086 0.161 0.136 0.083
DBW [36] 3.61 7.33 6.19 2.09 26.11 23.84 20.25 28.72 0.950 0.915 0.892 0.960 0.074 0.136 0.132 0.042
Ours (Block-level) 2.47 2.15 2.32 1.78 27.94 27.98 24.92 29.95 0.959 0.925 0.906 0.963 0.072 0.129 0.122 0.047
Ours (Point-level) 1.29 1.72 0.94 1.07 41.18 36.80 36.07 39.51 0.992 0.973 0.977 0.989 0.014 0.070 0.038 0.021
Table 4: Quantitative results on ShapeNet [6]. We report the Chamfer distance and novel view synthesis results across four categories.
4 Experiments
We comprehensively evaluate our approach across four key aspects: 3D reconstruction quality, view synthesis performance, shape parsimony, and computational efficiency, comparing it with state-of-the-art methods in the field.
4.1 Datasets
Extensive experiments are conducted on two widely-used public datasets: DTU [23] and ShapeNet [6]. Specifically, 15 standard scenes from the DTU dataset that are widely adopted for reconstruction quality assessment are used for evaluation. Each scene contains either 49 or 64 images. Additionally, we construct a ShapeNet subset consisting of four categories: Chair, Table, Gun, and Airplane, with 15 objects selected per category. For each object, 50 rendered images are generated for training. We further validate real-world applicability on both the BlendedMVS dataset [63] and self-captured scenes, demonstrating robust performance across synthetic and natural environments.
4.2 Implementation Details
We build upon the 2DGS [21] architecture and add a non-learnable part attribute to each Gaussian component to enforce the constraint in Eq. 13. In the ShapeNet benchmark, we observe a common degradation in mesh reconstruction as shown in Fig. 7. To address this, we propose Gaussian scale regularization, as validated in Sec. 4.4. We set the initial number of primitives $M$ to 8. The temperature parameter $\gamma$ in Eq. 8 is 0.005 and the overlapping number $k$ in Eq. 9 is 1.95. We train the hybrid representation for 30k iterations, followed by refinement optimization for another 30k iterations. During block-level optimization, the block-adding operation is carried out at the 5k-th and 10k-th iterations. The loss weights $\lambda_{\text{cov}}$, $\lambda_{\text{over}}$, $\lambda_{\text{par}}$, and $\lambda_{\text{opa}}$ are set to 10, 1, 0.002, and 0.01, respectively. For more details about datasets, metrics, and baselines, please refer to Supplementary.
4.3 Evaluations
4.3.1 DTU and Shapenet Benchmark
Geometry Reconstruction. In Tab. 1 and Tab. 3, we compare our geometry reconstruction to SOTA shape decomposition methods on Chamfer distance and training time using DTU and Shapenet dataset. The Chamfer distance metrics reported for ShapeNet are scaled by a factor of 100 for readability. Our method consistently outperforms in all scenes compared to prior works. As shown in Fig. 3, our approach consistently produces interpretable 3D decompositions with further refinement achieving more detailed geometry (the last two columns). MBF [42] achieves a low CD error but at the cost of using significantly more primitives, leading to over-decomposition and the inclusion of meaningless parts. Moreover, as shown in Tab. 2 and Tab. 4, our approach achieves competitive results with the advanced Gaussian-based 2DGS [21] and surface reconstruction Neuralangelo [30], which are limited to producing unstructured meshes. Notably, our model demonstrates excellent efficiency, with a reconstruction speed approximately 10X faster than PartNeRF [48] and over 3X faster than DBW [36].
Appearance Reconstruction. In addition to part decomposition, our method enables high-fidelity image synthesis, attributed to integrating Gaussian splatting within hybrid representations. EMS [32] and MBF [42] operate directly on point clouds and consequently lack image rendering capabilities, while PartNeRF [48] and DBW [36] yield low-quality view synthesis. In contrast, our method achieves high-quality appearance rendering results, as demonstrated in Tab. 2 and Tab. 4.
Figure 6: Qualitative comparison with SAM-based methods. Visual results demonstrate that our method produces more structurally coherent decompositions, whereas SAM-based approaches frequently exhibit spatial discontinuities.
Block-level Point-level
[…]
Complete model $2.32$ $24.92$ $0.906$ $0.122$ $\underline{0.94}$ $\underline{36.07}$ $0.977$ $0.038$ $6.6$
w/o $L_{\text{over}}$ $\underline{2.11}$ $\underline{25.17}$ $\underline{0.914}$ $\underline{0.118}$ $1.34$ $36.04$ $\underline{0.977}$ $\underline{0.037}$ $7.6$
w/o $L_{\text{opa}}$ $5.72$ $21.12$ $0.853$ $0.174$ $1.06$ $35.60$ $0.975$ $0.040$ $3.8$
w/o $L_{\text{cov}}$ $3.21$ $22.96$ $0.896$ $0.134$ $1.34$ $36.00$ $0.976$ $0.038$ $8.53$
w/o Adaptive $3.16$ $22.74$ $0.880$ $0.141$ $1.05$ $35.74$ $0.974$ $0.040$ $6.7$
w/o $L_{\text{par}}$ 1.91 25.48 0.918 0.115 0.93 36.18 0.977 0.037 $10.1$
Table 5: Ablation studies on the ShapeNet [6]. We report Chamfer Distance, rendering metrics, and the number of primitives (#P).
4.3.2 Real-life Data
To further demonstrate the applicability of our method for learning shape decomposition, we test our model on the real-life images from the BlendedMVS dataset [63] and a self-captured dataset. As shown in Fig. 12, our approach can robustly produce both realistic appearances and reasonable 3D decompositions across a variety of data types. More results are provided in the supplementary material.
Figure 7: Impact of Gaussian scale regularization (SR). The degraded mesh produced by 2DGS [21] is effectively improved.
Figure 8: Impact of the initial number of primitives ($M$). Increasing $M$ yields a finer-grained decomposition, while decreasing it produces a coarser decomposition.
Figure 9: Applications. Part-Aware Editing (top): After optimization, we can easily edit the scene by adding, scaling, or moving specific parts. 3D Content Generation (bottom): By combining parts of different objects, we can create new 3D content.
4.3.3 Compared to SAM-based Methods
We also conduct comparisons to SAM-based [27] methods for this task, as shown in Fig. 6. Despite impressive advances in 3D segmentation and editing [5, 26, 65, 66] by distilling 2D information, achieving semantic disentanglement but remains challenging due to indistinct textures and severe cross-view 2D inconsistencies. Lifting 2D segmentation features to 3D segmentation will also lead to feature misalignment, leading to visual artifacts. In contrast, our method performs direct 3D space segmentation, yielding more reasonable and structurally coherent decompositions.
4.4 Ablations
Block-level Point-level
[…]
$\lambda_{\text{par}}$ = 0.1 $6.46$ $16.10$ $0.721$ $0.255$ $1.08$ $34.30$ $0.987$ $0.018$ $2.7$
$\lambda_{\text{par}}$ = 0.01 (*) $4.19$ $19.84$ $0.820$ $0.189$ $1.05$ $35.04$ $0.988$ $0.015$ $5.9$
$\lambda_{\text{par}}$ = 0.001 $4.01$ $20.18$ $0.835$ $0.180$ $1.03$ $35.04$ $0.988$ $0.015$ $7.1$
$M$= 4 $4.97$ $19.06$ $0.794$ $0.213$ $1.06$ $34.93$ $0.987$ $0.016$ $5.0$
$M$= 8 (*) $4.19$ $19.84$ $0.820$ $0.189$ $1.05$ $35.04$ $0.988$ $0.015$ $5.9$
$M$= 16 $3.99$ $20.87$ $0.83$ $0.176$ $1.07$ $35.05$ $0.988$ $0.015$ $8.3$
Table 6: Effect of parsimony weight ($\lambda_{\text{par}}$) and initial primitives count ($M$) on DTU [23]. We report Chamfer distance, rendering metrics, and primitives count (#P). * denotes the default setting.
In this section, we first analyze the design choices by isolating each strategy to assess its impact. We report the averaged quantitative performance on 15 instances of the chair category in Tab. 5 and provide visual comparisons in Fig. 5. Removing overlapping loss provides better reconstruction accuracy but more overlapping parts. The opacity loss drastically affects the number of primitives and doubles the CD, highlighting its role in controlling primitive presence and reconstruction quality. Coverage loss is crucial in reconstruction accuracy, ensuring primitives align correctly with the target. The adaptive primitive strategy fills in missing parts and improves the reconstruction metric. Without parsimony loss, CD and rendering metrics are optimal but result in over-decomposition. Note that the full model did not yield the highest accuracy because the part-aware reconstruction trades off the accuracy and rationality of the part decomposition. The objective is to achieve high accuracy while preserving reasonable parts.
We also validate the effectiveness of Gaussian scale regularization in Fig. 7. 2DGS [21] tends to use large-scale Gaussians in texture-less areas and create holes in the constructed mesh by TSDF integration [71]. The presence of holes stems from the inability of large-scale Gaussians to maintain view-consistent depth. Scale regularization effectively addresses this issue by suppressing large-scale Gaussians.
Lastly, in Tab. 6, we analyze the influence of two key hyperparameters on decomposition granularity: the parsimony loss weight $\lambda_{\text{par}}$ and the initial primitive count $M$. Stronger parsimony regularization reduces the number of primitives, while weaker regularization increases it. As $M$ rises, both reconstruction and view synthesis performance slightly improve. This demonstrates that adjusting $M$ and $\lambda_{\text{par}}$ enables our method to effectively control the granularity of object or scene decomposition. Fig. 8 visually illustrates this impact.
4.5 Applications
Fig. 9 illustrates two applications of our method that the original 3DGS and NeRF-based approaches do not support. Firstly, after optimization, we obtain the part decomposition, facilitating easy editing of specific objects or scene components, e.g., adding, moving, removing, or scaling. Secondly, by combining parts of different objects, our method enables the creation of new high-quality 3D content.
5 Conclusions
We introduce PartGS, a hybrid representation of superquadrics and 2D Gaussians, to learn 3D scenes in a part-aware representation. Compared to prior works, the proposed method retains geometry details, supporting high-quality image rendering. It obtains state-of-the-art performance in comprehensive evaluations. One limitation is that it takes background-free scene images as inputs, leveraging segmentation tools such as SAM [27]. In the future, we aim to explore how to model backgrounds and extend to larger and more complex scenes.
Acknowledgements
This work is supported in part by the NSFC (62325211, 62132021, 62372457), the Major Program of Xiangjiang Laboratory (23XJ01009), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), the Natural Science Foundation of Hunan Province of China (2022RC1104).
Supplementary Material
At first, this supplementary material provides detailed experimental settings, including data processing procedures, implementation details for comparison baseline methods, and evaluation metrics (Sec.6). Subsequently, we provide further implementation specifics, including the training configurations and rendering process (Sec.8).
In Fig. 10 and Fig. 11, we provide additional visual comparisons of our method against state-of-the-art baselines on DTU [23] and ShapeNet [6] datasets. Moreover, we include visual results on several scenes from the BlendedMVS [63] and self-captured data in Fig. 12. It can be observed that the decomposed reconstructions by our method are more reasonable and precise, which is believed to be beneficial for downstream tasks. Our code and data are available at https://github.com/zhirui-gao/PartGS.
6 More Details on Experiments
6.1 Datasets
We conduct evaluations on two public datasets: DTU [23] and ShapeNet [6]. DTU is a multi-view stereo (MVS) dataset comprising 80 forward-facing scenes, each captured with 49 to 64 images. In experiments, we use 15 publicly recognized scenes commonly adopted in previous studies [21, 30]. We downsample the images to a resolution of $400\times 300$ for computational efficiency.
To validate the effectiveness of our method in the decomposition of man-made objects, we construct a subset of the ShapeNet dataset comprising four categories: Chair, Table, Gun, and Airplane. Each category includes 15 distinct objects, providing diverse instances. For each object, we randomly sample 100 camera poses on a sphere and render images at a resolution of $400\times 400$. The rendered images are equally split into training and testing sets.
Additionally, to explore the potential of PartGS in handling real-life data, we present qualitative results on the BlendedMVS dataset [63] and self-captured scenes. For BlendedMVS, we use official camera poses, while for self-captured scenes, camera poses are estimated by COLMAP [44] and normalized using IDR [64]. All images are resized to $400\times 300$ pixels. For the DTU, ShapeNet, and BlendedMVS datasets, we utilize the ground-truth foreground masks provided within the datasets. For self-captured real-world scenes, we employ the Segment Anything Model (SAM) [27] to segment foreground objects.
6.2 Implementation Details on Baselines
We compare our method with four state-of-the-art works on 3D shape decomposition: EMS [32], MonteboxFinder (MBF) [42], PartNeRF [48] and DBW [36]. MBF and EMS are applied to point clouds, utilizing cuboids and superquadrics, respectively, to fit the 3D points. The input point clouds are sampled from either ground truth 3D shapes (GT) or meshes reconstructed by state-of-the-art MVS method Neus [55]. We adhere to the testing procedure outlined by DBW [36] and randomly sample 5K and 200K points from the GT meshes as the input for EMS and MBF, respectively. Similar to our problem setting, PartNeRF and DBW use images as input to construct structural 3D representations. To ensure fair comparisons, we discard the pre-defined ground plane in DBW, as the test scenario consists exclusively of foreground objects. For PartNeRF, we perform instance-specific training and set the same number of parts as in our method, while retaining all other default parameter settings. Additionally, we compare with SOTA Gaussian Splatting reconstruction method 2DGS [21], and surface reconstruction approach Neuralangelo [30], to provide an intuitive evaluation of the reconstruction quality achieved by our method.
6.3 Evaluation Metrics
We evaluate from 3D reconstruction quality, view synthesis, shape parsimony, and reconstruction time.
  • The 3D reconstruction quality is measured by the official Chamfer distance (CD) evaluation [23] between the recovered geometry and GT, reflecting the accuracy of 3D reconstructions.
  • View synthesis uses three standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [57], and Learned Perceptual Image Patch Similarity (LPIPS) [69].
  • Shape parsimony is quantified by the number of parts, while reconstruction time refers to the running time measured on the same device. While shape parsimony does not directly indicate the quality of the decomposition, a smaller number of decomposed components, given consistent reconstruction quality, suggests a more concise and reasonable decomposition. This is because fitting different parts of an object with fewer blocks is inherently more challenging, whereas utilizing more blocks simplifies the fitting process.
Figure 10: Qualitative comparisons on DTU [23]. We compare our approach with state-of-the-art baselines on the DTU dataset with the background removed. The last two columns show our block-level and point-level reconstructions, respectively. Our method is the only one that provides reasonable 3D part decomposition while capturing detailed geometry.
Figure 11: Qualitative comparisons on ShapeNet [23]. We compared our approach with state-of-the-art baselines across four categories. The last two columns display our block-level and point-level reconstructions, respectively. Our method uniquely provides reasonable 3D part decomposition while simultaneously capturing detailed geometry.
Figure 12: Qualitative results on BlendedMVS [63] and self-capatured data. We demonstrate the RGB renderings and decomposed parts from novel views across a variety of objects. The first five examples are from the BlendedMVS dataset, and the remaining examples are from our own captured scenes.
7 Additional Comparative Experiments
Very recently, DPA-Net [68] and GaussianBlock [24] have achieved advanced 3D shape abstraction. DPA-Net enables part-aware reconstruction in a feedforward manner, requiring approximately 3 days of GPU training on a pre-collected dataset and 2 hours per object for inference. GaussianBlock utilizes SAM [27] to guide superquadric splitting and fusion for 3D decomposition, with a processing time of 6 hours per object. We compare our method on DTU and BlendedMVS scenes, with results presented in Fig. 13. The proposed approach achieves comparable performance while demonstrating superior efficiency. It is important to clarify the supervision requirements: DPA-Net is a supervised method that relies on SAM-generated segmentation masks for training, whereas GaussianBlock, while not strictly 3D supervised, depends on high-quality pre-trained datasets that typically require days of training. Our method distinguishes itself by being completely self-supervised, requiring neither 3D supervision nor pre-trained components, yet achieving significantly better computational efficiency.
Figure 13: Qualitative comparisons to DPA-Net and GaussianBlock. The first two examples are from the DTU dataset, and the last two examples are from the BlendedMVS dataset.
8 Implementation Details
8.1 Training Configurations
To ensure uniform distributions of 2D Gaussians across surfaces of the mesh, random barycentric coordinates are generated directly within each triangular face. Specifically, barycentric weights are computed as $u=\sqrt{rand}$ and $v=rand$. These weights are transformed to obtain $\alpha=[1-u,u(1-v),uv]$, $rand\in(0,1)$. The sampled position on triangular face is calculated as $o=\alpha_{0}v0+\alpha_{1}v1+\alpha_{2}v2$, where $v_{o},v_{1},v_{2}$ are vertices of the triangle. This method compensates for the non-uniformity caused by the triangle’s geometry. Unlike naive random sampling which results in uneven distributions, it ensures that the sampled points are evenly distributed across mesh surfaces.
The same hyperparameters are used for all experiments. We set the initial number of primitives $M$ to 8. In the hybrid representation, each superquadric mesh is a level-2 icosphere (320 triangular face). Each triangular face contains 100 Gaussians, with a scaling parameter $c$ of 0.1. The number of sampled points in each ray is 2048.
During refinement, we employ regularization in 2DGS to achieve better geometric reconstruction, including depth distortion maps, depth maps, and normal maps. Additionally, we introduce a mask cross-entropy to filter out extra noise Gaussians. To extract the meshes from 2D Gaussians, we use truncated signed distance fusion (TSDF) to fuse rendered depth maps, utilizing Open3D [71].
8.2 Rendering
With Gaussians attached to the surface of each block, we achieve view-dependent rendering through tile-based rasterization as Gaussian Splatting [25, 21]. Given a view, Gaussians on the superquadric surface are projected onto the image space, forming an RGB image. Initially, the screen space determines the bounding box for each Gaussian. Subsequently, these Gaussian ellipses are sorted according to their depths of center to the image plane. Finally, volumetric alpha compositing [41] is utilized to integrate the alpha-weighted RGB values for each pixel.
To formulate the process, considering a viewing transformation ${W}$, the covariance matrix ${\Sigma}_{i}^{{}^{\prime}}$ of $i$-th Gaussian in the camera coordinate system is calculated by:
${\Sigma}_{i}^{{}^{\prime}}={J}{W}{\Sigma}{W}^{T}{J}^{T},$ (14)
where ${J}$ is the Jacobian of the affine approximation of the projective transformation, and ${\Sigma}$ is the covariance matrix of the Gaussian ellipse. Note that the last row and column of ${\Sigma}$ are omitted since we adopt 2D Gaussians. Following alpha compositing, we first calculate an alpha value for each Gaussian ellipse:
$\alpha_{i}=\tau_{i}\exp(-\frac{1}{2}(\mathbf{x}-\mathbf{u}_{i})^{T}{\Sigma_{i}^{{}^{\prime}}}_{-1}(\mathbf{x}-\mathbf{u}_{i}))).$ (15)
Here, $\mathbf{u}_{i}$ is the center coordinate of the projected Gaussian ellipse, and $\tau_{i}$ is the opacity of the block where the $i$-th Gaussian is located. The calculated alphas are sorted according to their depths from the image plane. Meanwhile, we can acquire the color value $c_{i}$ from the spherical harmonics of $\mathcal{N}$ ordered points, thus obtaining the rendering RGB value:
$C=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (16)
This process is differentiable and can optimize the hybrid representation through gradient descent.